OpenStudy (anonymous):

The volume of water remaining in a hot tub when it is being drained satisfies the differential equation dV/dt = −2 ( V)^(1/2) , where V is the number of cubic feet of water that remain t minutes after the drain is opened. Find V if the tub initially contained 121 cubic feet of water.

OpenStudy (anonymous):

First separate variables and solve the differential equation for V

OpenStudy (anonymous):

\[dV/V ^{1/2} = -2dt\]

OpenStudy (anonymous):

so \[\int\limits_{?}^{?}dV/V ^{1/2} = \int\limits_{?}^{?} -2dt\]

OpenStudy (anonymous):

where both of those are indefinite integrals

OpenStudy (anonymous):

alright great! I've got it from there! Thank you very much.

OpenStudy (anonymous):

No problem, let me know if you get stuck anywhere else with this one

OpenStudy (anonymous):

\[V = t ^{2}-22t+121\] was the final answer i found

OpenStudy (anonymous):

yes i did too except i left it as (-t+11)^2

OpenStudy (anonymous):

perfect :)

OpenStudy (anonymous):

thanks again!